### v1

# A Linear Kernel for Planar Vector Domination

Abstract
Given a graph $G$, an integer $k\geq 0$, and a non-negative integral function $f:V(G)\rightarrow\mathcal{N}$, the Vector Domination problem asks whether a set $S$ of vertices, of cardinality $k$ or less, exists in $G$ so that every vertex $v\in V(G)-S$ has at least $f(v)$ neighbors in $S$. The problem generalizes several domination problems and it has also been shown to generalize Bounded-Degree Vertex Deletion. In this paper, the parameterized version of Vector Domination is studied when the input graph is planar. A linear problem kernel is presented.
Keywords:
Dominating set Vector dominationBounded-degree vertex deletion Parameterized Complexity Kernelization
1 Introduction
The Vector Dominating Set problem assumes a graph $G$ is given along with an integer $k\geq 0$ and a demand function $f$ from the set $V$ of vertices of $G$ to the set $\{0,\ldots|V|-1\}$, and asks whether a set $D$ of at most $k$ vertices of $G$ exists such that each vertex $v$ has at least $f(v)$ neighbors in $D$. The problem can be viewed as a generalization of many domination problems that are extensively studied in the literature, including the classic Dominating Set, the $k$-dominating set problem [18, 4, 5, 6, 9, 15] and the bounded degree deletion problem [12, 3, 10] along with its dual problem the $s$-Plex Detection problem, which finds application in social network analysis [20, 2].
From a computational complexity standpoint, Vector Dominating Set is $NP$-hard being a generalization of the dominating set problem. Thus it is natural to ask under what conditions the problem becomes tractable. It is well known by now that the study of a problem’s parameterized complexity [7, 8] provides an alternative view of its complexity and can shed light on classes/types of input where the problem can be tractable. Under the parameterized complexity lens, a problem is fixed parameter tractable (FPT) if there exists an algorithm that solves it in $O(f(k)n^{c})$ time where $n$ is the size of the input, $k$ is an input parameter, $f$ is a function of $k$ and $c$ is a constant. Alternatively, a problem is FPT if it admits a kernel i.e given an instance $(G,k)$ of the problem, an equivalent instance $(G^{\prime},k^{\prime})$ can be constructed in polynomial time where $|G^{\prime}|\leq g(k)$ and $k^{\prime}\leq k$. $G^{\prime}$ is referred to as a problem kernel. Of particular interest is the case where $g$ is a (low degree) polynomial function of $k$. The dominating set problem is $W[2]$-hard on general graphs [11], thus there is little or no hope in achieving fixed parameter tractability for it or for the more general vector dominating set problem on general graphs.
Although planar vector domination is known to be fixed-parameter tractable [17], a natural question to ask is whether the problem admits a polynomial kernel on classes of graphs where it is known to be fixed-parameter tractable. On planar graphs, Alber et al. (2004) provided an explicit linear kernel for the dominating set problem. This was based on structural properties of planar graphs, especially degree structure, and it was followed by a series of papers on linear kernels of different domination problems [1, 14, 13, 19]. No linear kernel is known yet for Vector Dominating Set in planar graphs. Addressing this problem is the main objective of this paper. A linear kernel is presented based on several structural properties of planar graphs and the approach of Alber et al.. The bound achieved is $101k$, which is comparable to the upper bounds for related problems. Moreover, the obtained bound leads to a linear kernel for the Planar Bounded-Degree Vertex Deletion problem.
2 Preliminaries
Common graph-theoretic terminology is used in this paper, with a main focus on planar graphs and their properties. Although we assume the input to the considered problem is a planar graph that is given with a fixed embedding (i.e., plane graph), we note that deciding whether a given graph is planar, and constructing a corresponding planar embedding, can be done in linear time [16]. It is well known that the number of edges in a planar graph $G$ with $|G|\geq 3$ is at most $63|G|-$ 6, where $|G|$ is the order of $G$ (i.e., number of vertices). Moreover, if $G$ is bipartite the number of edges is at most $42|G|-$ 4. The following problem is considered:
Planar Vector Dominating Set (PVDS)
Given: A planar graph $G=(V,E)$ of order $n$, an $n$-dimensional non-negative demand vector $d$, and an integer $k$.
Question: Is there a set $S\subset V$ of cardinality $k$ such that every vertex $v\in G-S$ has at least d(v) neighbors in $S$?
Let $(G,d,k)$ be an instance of PVDS. A vertex $v\in V(G)$ is said to be a $j$-vertex if $d(v)=j$ where $j\in\mathbb{N}$. A $j$-vertex will be considered demanding, or of high-demand if $j>0$, otherwise it will be a low-demand vertex. We denote by $G-v$ the subgraph of $G$ induced by $V(G)\setminus\{v\}$. The set of neighbors of $v$ is denoted by $N(v)$ while $N_{l}(v)$ and $N_{h}(v)$ are sets of low-demand and high-demand neighbors of $v$, respectively. Furthermore, we denote by $N[v]$ the set $N(v)\cup\{v\}$. The same applies to $N_{l}[v]$ and $N_{h}[v]$.
For a set of vertices $A\subset V(G)$ we define $N(A)=\cup_{v\in A}N(v)$, and we denote by $N_{l}(A)$ and $N_{h}(A)$ the sets of low-degree and high-degree elements of $N(A)$, respectively. $N[A],N_{l}[A]$ and $N_{h}[A]$ are defined analogously.
A solution is said to observe an edge $e=uv$ where $u,v\in V(G)$ if one of the endpoints of $e$, namely $u$, is deleted to decrease d(v). We say that a set $A$ dominates another set $B$ if for all $v\in B$, $|N(v)\cap A|\geq d(v)$.
In some cases, we know certain vertices are not to be selected into a PVDS solution. This gives rise to a yet another more general problem. In particular, we shall consider the following annotated PVDS problem:
Annotated Planar Vector Dominating Set (APVDS)
Given: A planar graph $G=(V,E)$ of order $n$, an $n$-dimensional non-negative demand vector $d$, and an integer $k$, and a set $P\subset V$.
Question: Is there a set $S\subset V-P$ of cardinality at most $k$ such that every vertex $v\in G-S$ has at least d(v) neighbors in $S$?
Redundant solution elements play a role in our kernelization method. Let $v$ be a vertex such that: if there is a solution $S$ of APVDS where $v\in S$, then there is anther solution $S^{\prime}$ where $v\notin S^{\prime}$. We will further annotate $G$ by assigning a special blue color to such vertices, which are assumed to be redundant (or replaceable) in this case. Note that if $v$ is redundant, then the two APVDS instances $(G,k,d,P)$ and $(G,k,d,P\cup\{v\})$ are equivalent.
3 A Linear Kernel
The following reduction rules for the PVDS are assumed to be applied successively and exhaustively. The first five rules are applicable to the general Vector Dominating Set problem.
Reduction Rule 1
If $u$ and $v$ are adjacent 0-vertices then delete the edge $uv$.
Reduction Rule 2
Delete all isolated 0-vertices.
Reduction Rule 3
If for some vertex $v$, $d(v)>k$ or $d(v)>|N(v)|$, delete $v$, then decrease $d(u)$ by one for all $u\in N(v)$ and decrease $k$ by one.
Reduction Rule 4
Let $v$ be a 0-vertex. If there exists a vertex $a\in V(G)$ such that $N(v)\subset N[a]$, then for any 1-vertex $b$ in $N(v)$, delete the edge $vb$. Delete the edge $va$ if it exists (see Figure 1).
xyavztb
Figure 1: Reduction Rule 4: delete $vb$.
Soundness. Any potential solution must contain a neighbor of vertex $b$. If this neighbor is $v$, then it can be very well replaced by $a$. Therefore, if a solution exists, we know for sure there must be a solution that does not observe the edges $bv$ and $av$.
Reduction Rule 5
Let $v$ be a 1-vertex. If there exists a vertex $a\in N(v)$ such that every vertex $u\in N[v]-a$ satisfies: $d(u)\leq 1$, and $N_{h}[u]\subset N[a]$, then delete $a$, decrease the demand of each neighbor of $a$ by one, and decrease k by one.
Soundness. Any solution $S$ must contain a vertex $b$ of $N[v]$. If $a\notin S$, replace $b$ by $a$ since $N_{h}[b]\subset N[a]$.
Definition 1
For $s_{1},s_{2}\in V(G)$, $s_{1}\neq s_{2}$:
  • An $s_{1}s_{2}-$path of length two is called a type 1 path.
  • An $s_{1}s_{2}-$path of length 4 $s_{1}vcv^{\prime}s_{2}$ is called a type 2 path if $d(c)=0$, $d(v)=1$, $v\notin N(s_{2})$ and $v^{\prime}\notin N(s_{1})$.
  • An $s_{1}s_{2}-$path of length 3 $s_{1}vv^{\prime}s_{2}$ is called a type 3 path if $d(v)\leq 1$. A 1-vertex outside this path adjacent to both $s_{1}$ and $v$ is called a "leech" for this path.
Definition 2
Suppose the graph contains a solution set $S$. A real path $p$ is a path with endpoints in $S$ satisfying one of the following conditions:
  • $p$ is a type 1 path;
  • $p$ is a type 2 path such that $p=s_{1}vcv^{\prime}s_{2}$ where $v$ and $v^{\prime}$ are not both used by a type 1 real path;
  • $p$ is a type 3 path such that $p=s_{1}vv^{\prime}s_{2}$ where $v$ and $v^{\prime}$ are not both used by a type 1 real path.
From this point one, we denote by $A$ the set of 0-vertices such that $N(A)\subset\{u\in V(G):|N(u)\cap S|\geq 2\}$. Set $L$ to be the set of all leeches for real paths.
Lemma 1
Suppose V(G) contains a solution set $S$, then after applying the reduction rules above any vertex $v\in G-(S\cup A\cup L)$ belongs to some real path.
Proof
Suppose V(G) contains a deletion set $S$. Let $v\in G\setminus(S\cup A\cup L)$. If $d(v)\geq 2$, $v$ must be adjacent to two vertices in $S$ so $v$ is on a real path of type 1. If $d(v)=1$, it must be adjacent to some $s_{1}\in S$. We may assume that $N(v)\cap S=\{s_{1}\}$ otherwise it lies on a type 1 real path. If it is connected to a 0-vertex $c$, $c\notin S$, due to reduction rule 4 we have $N(c)\not\subset N[s_{1}]$. Let $v^{\prime}\in N(c)\setminus N[s_{1}]$, Due to reduction rule 1 $v^{\prime}\in N[s_{2}]$ for some vertex $s_{2}\in S$, $s_{2}\neq s_{1}$. We may easily verify that $v$ is on some real path. Now let $u\in N(v)\setminus S$. If $|N(u)\cap S|\geq 2$, $v$ lies on a type 3 real path, so we may assume that $d(u)=1$ for all $u\in N(v)-s_{1}$. Due to reduction rule 5, there exists a vertex $u^{\prime}\in N(v)-s_{1}$ where $u^{\prime}\in N(s_{2})$ for some $s_{2}\in S,s_{2}\neq s_{1}$. Thus $v$ is a leech for the real path $s_{1}uu^{\prime}s_{2}$, a contradiction.
Now suppose that $d(v)=0$. If $|N(v)\cap S|\geq 2$, $v$ lies on a type 1 real path, and as above if $N(v)\cap S=\{s_{1}\}$ then there exists a real path $s_{1}vus_{2}$. So we may suppose that $N(v)\cap S=\emptyset$. Since $v\notin A$, there exists $u\in N(v)$ where $N(u)\cap S=\{s_{1}\}$, As above using reduction rule 4 we may verify that there exists a real path $s_{1}uvu^{\prime}s_{2}$. ∎
Following the approach of Alber et al (2004), we define the concept of a region.
Definition 3
Let $G(V,E)$ be a plane graph with solution set $S$. A region $R(v,w)$ between two vertices $v,w\in S$ is a closed subset of the plane with the following properties:
  • The boundary of $R(v,w)$ is formed by two $vw$-real paths $o_{1}$ and $o_{2}$;
  • All vertices strictly inside the region $R(v,w)$ can be dominated by $\{v,w\}$;
  • There is a line in the plane separating the region $R(v,w)$ into two closed areas $A_{1}$ and $A_{2}$ where $o_{1}$ lies on the boundary of $A_{1}$ and $o_{2}$ lies on the boundary of $A_{2}$.
From this point on, we refer to the above paths $o_{1}$ and $o_{2}$ as outer paths between $v$ and $w$. For a region $R=R(v,w)$, let $V(R)$ denote the vertices belonging to $R$, i.e., $V(R)=\{u\in V:u$ sits inside or on the boundary of $R\}$. In what follows, the boundary of a region $R$ will be denoted by $\partial R$.
Definition 4
[1] Let $G=(V,E)$ be a plane graph and $S\subset V$. An S-region decomposition of $G$ is a set $\mathscr{R}$ of regions between pairs of vertices in $S$ such that:
  1. 1.
[…]
  2. 2.
    for two regions $R_{1},R_{2}\in$$\mathscr{R}$, $(R_{1}\cap R_{2})\subset(\partial R_{1}\cup\partial R_{2})$.
An S-region decomposition $\mathscr{R}$ is called maximal if it is maximal for inclusion.
One defines the graph $G_{\mathscr{R}}=(V_{\mathscr{R}},E_{\mathscr{R}})$ with possible multiple edges of an S-region decomposition $\mathscr{R}$ of $G$ where $V_{\mathscr{R}}=S$ and $E_{\mathscr{R}}=\{\{v,w\}:$ there is a region $R(v,w)\in\mathscr{R}$ between $v,w\in S\}$.
We shall show that after applying the reduction rules above to an instance $(G,k)$ of PVDS, for which a solution set $S$ exists, there exists a maximal $S$-region decomposition $\mathscr{R}$ such that:
  1. 1.
    The number of regions of $\mathscr{R}$ is $O(k)$.
  2. 2.
    Each region $R\in$$\mathscr{R}$ contains $O(1)$ vertices.
  3. 3.
    The number of vertices not in $V(\mathscr{R})$ is $O(k)$.
These facts together mean that the size of the graph after applying the reduction rules is $O(k)$, and a linear kernel is obtained.
Definition 5
[1] A planar graph $G=(V,E)$ with multiple edges is thin if there exists a planar embedding such that if there are two edges $e_{1}$, $e_{2}$ between a pair of distinct vertices $v,w\in V$, then there must be two further vertices $u_{1},u_{2}\in V$ which sit inside the two disjoint areas of the plane that are enclosed by $e_{1}$, $e_{2}$.
Lemma 2
[1] For a thin planar graph G = (V, E) we have $|E|\leq 3|V|$-$6$.
The following proposition can be easily deduced from [1].
Proposition 1
For a plane graph G with vector dominating set S, there exists a maximal S-region decomposition $\mathscr{R}$ such that $G_{\mathscr{R}}$ is thin.
Consequently, the number of regions $R(v,w)$ will be bounded by $63k-$ 6.
Lemma 3
Suppose $G$ admits a solution $S$ with a maximal S-region decomposition $\mathscr{R}$, then any vertex $v\in G-(S\cup A\cup L)$ is in $V(\mathscr{R})$.
Proof
Take $v\notin L\cup A\cup S$, then by Lemma 1 $v$ must belong to some real path $p$. We will say that an edge crosses a region $R$ if the edge lies (except possibly for its endpoints) strictly inside $R$. Similarly, we say that a path crosses a region $R$ if at least one edge of the path crosses $R$. Suppose that $v\notin V(\mathscr{R})$, then $p$ must cross some region $R=R(s_{1},s_{2})$ where $s_{1},s_{2}\in S$. Let $o_{1}$ and $o_{2}$ be the two outer paths of $R$, then $p$ must intersect $o_{1}$ or $o_{2}$. Without loss of generality, we will assume $p$ intersects $o_{1}$. If $p$ is a type 1 path the Lemma follows trivially. Thus we will consider the following cases:
  • Case 1: $p$ is a type 2 path $svcv^{\prime}s^{\prime}$.
    In this case either $c\in o_{1}$ or $v^{\prime}\in o_{1}$. If $c\in o_{1}$ and $o_{1}$ is a type 1 path a real path disjoint from $\mathscr{R}$ arises, a contradiction.
    If $o_{1}$ is a type 2 path, $o_{1}=s_{1}ucu^{\prime}s_{2}$, then one of the regions $svcus_{1}$ or $svcu^{\prime}s_{2}$ contains $v$ contradicting the maximality of $\mathscr{R}$. Indeed we may suppose that $s\neq s_{2}$, if $u^{\prime}$ and $s$ are not adjacent, then $svcu^{\prime}s_{2}$ is a real path disjoint from $\mathscr{R}$. Otherwise $s\neq s_{1}$ then $svcus_{1}$ has the desired property as $u$ and $s$ are not adjacent due to the fact that $o_{1}$ is real. Now if $o_{1}$ is a type 3 path $s_{1}cus_{2}$ then if $s\neq s_{1}$ the path $svcs_{1}$ is a region that contains $v$, else $s=s_{1}$ then $v$ is a leech for $o_{1}$ so $v\in L$ contradicting our assumption. Now if $c\notin o_{1}$ and $v^{\prime}\in o_{1}$. In this case the edge $v^{\prime}s^{\prime}$ crosses $R$ so $s^{\prime}\in V(R)$ thus $s^{\prime}$ is an end of $o_{1}$. Then $o_{1}$ is not of type 1 or 2, since in these cases $v^{\prime}s^{\prime}\in o_{1}$. The remaining case where $o_{1}$ is a type 3 path may be treated exactly as above.
  • Case 2: p is a type 2 path $suvu^{\prime}s^{\prime}$ or a type 3 path $svu^{\prime}s^{\prime}$.
    In this case suppose (namely) that the edge $u^{\prime}s^{\prime}$ crosses $R$, then, as in the previous argument, a region containing $u$ and $v$ can be found that does not cross any other region in $\mathscr{R}$, contradicting the maximality of $\mathscr{R}$, or the same reasoning as above can be repeated.
Definition 6
let $a_{1},a_{2}\in V(G)$, Call a closed subset of the plane $C(a_{1},a_{2})$ a candidate region between $a_{1}$ and $a_{2}$ if the boundary of $C(a_{1},a_{2})$ is formed of two $a_{1}a_{2}-$paths of type 1,2 or 3, every vertex in the interior of $C(a_{1},a_{2})$ is dominated by $\{a_{1},a_{2}\}$, and $C(a_{1},a_{2})$ is maximal for these properties (for inclusion).
Let $a_{1},a_{2}\in V(G)$ and take $C(a_{1},a_{2})$ to be a candidate region between $a_{1}$ and $a_{2}$ with boundary $O(a_{1},a_{2})=O_{1}\cup O_{2}$. Consider the following sets:
  • $Y(a_{1},a_{2})=\{v:v\in O(a_{1},a_{2})$ and $d(v)\geq 2\}$
  • $B(a_{1},a_{2})$ is the set of vertices in the interior of $C(a_{1},a_{2})$ that are adjacent to some vertex in $O(a_{1},a_{2})$.
  • $I(a_{1},a_{2})$ is the set of vertices in the interior of $C(a_{1},a_{2})$ that are not in $B(a_{1},a_{2})$.
  • $I^{\prime}(a_{1},a_{2})=$$\{v$: $\{v\}$ dominates $I(a_{1},a_{2}),v\notin P\}$
  • $O^{\prime}(a_{1},a_{2})=\{v:|N(v)\cap Y(a_{1},a_{2})|\geq 2\}$
In the following rules let $a_{1},a_{2}\in V(G)$,$C(a_{1},a_{2})$ be a candidate region between $a_{1}$ and $a_{2}$. Let $C,O,B,I,I^{\prime}$ and $O^{\prime}$ denote $C(a_{1},a_{2})$, $O(a_{1},a_{2})$, $B(a_{1},a_{2}),$ $I(a_{1},a_{2}),$ $I^{\prime}(a_{1},a_{2})$ and $O^{\prime}(a_{1},a_{2})$ respectively. Reduction rules $6,7$ and $8$ apply only on planar graphs. They are not applied exhaustively.
Note that if $I\neq\emptyset$ and a vertex $w\notin I^{\prime}$ is in the solution set $S$, then there must be another vertex $w^{\prime}\in I\cup B\cup\{a_{1},a_{2}\}$ in $S$. The following reduction rule follows:
Reduction Rule 6
Suppose $I\neq\emptyset$ and let $u\in C$. If $u\notin I^{\prime}$ and $N(u)\cap Y=\emptyset$, then color $u$ blue.
Soundness. Suppose $u\in S$, since $u\notin I^{\prime}$, $\exists u^{\prime}\in I\cup B\cup\{a_{1},a_{2}\}$ in $S$. We have $Y\cap N(u)=\emptyset$, then if $u^{\prime}=a_{1}$ (namely) replace $u$ by $a_{2}$ (all $N[u]$ is dominated by $\{a_{1},a_{2}\}$). So suppose $u^{\prime}\in B\cup I$, then replace $u$ and $u^{\prime}$ by $a_{1}$ and $a_{2}$, since $B\cup I$ would be dominated and $|N(Y)\cap\{u,u^{\prime}\}|\leq 1$ while $\forall y\in Y,|N(y)\cap\{a1,a2\}|\geq 1$.
Reduction Rule 7
Suppose $I\neq\emptyset$ and $O^{\prime}\neq\emptyset$. If a vertex $w$ is inside $C$, then color it blue unless one of the following holds:
  • $w\in O^{\prime}\cup I^{\prime}$.
  • $\exists w_{2}\in N(Y)$ where $\{w,w_{2}\}$ dominates $I$.
  • $\exists w_{2}\in O^{\prime}$ where $\{w,w_{2}\}$ dominates $I\cap N(a_{1})$ or $I\cap N(a_{2})$.
Soundness. Suppose $w\notin O^{\prime}\cup I^{\prime}$ and $w\in S$. We will show that if $w$ is a potential solution vertex then one of the above cases must hold. $w$ is a potential solution vertex so by reduction rule 6 it must have a neighbor in $Y$. Let $y\in N(w)\cap Y$ with $y\in N(a_{1})$ (namely) and $y^{\prime}$ be the other vertex of $Y$. Let $W$ be the vertices in $S\cap C$ except for vertices in $O-Y$. We have $|W|\leq 3$ since any four vertices in $W$ can be replaced by $a_{1},a_{2},y$ and $y^{\prime}$. If the vertices of $W$ are all strictly inside $C$ with $W\cap O^{\prime}=\emptyset$ and $|W|\geq 3$, then we must have either $|N(y)\cap W|\leq 1$ or $|N(y^{\prime})\cap W|\leq 1$. Suppose namely it is the first case then the vertices of $W$ can be replaced by $a_{1},a_{2},$ and $y^{\prime}$. Now we may assume that $W$ contains vertices of $O\cup O^{\prime}\cup\{a_{1},a_{2}\}$. If $W\cap O\neq\emptyset$, $w$ and at most 1 other vertex in $W$ must dominate $I$ ($I\cap N(O)=\emptyset$) thus $w$ satisfies one of the cases of the reduction rule. Now if $W\cap O=\emptyset$ then if we take $w^{\prime},w^{\prime\prime}$ in $W$ different from $w$, it can be easily verified that any combination of choices of $w^{\prime},w^{\prime\prime}$ either leads to replacing a vertex of $W$ or to showing that $w$ follows one of the above cases. An example would be if $w^{\prime},w^{\prime\prime}\in O^{\prime}$ then if $Y\subset N(a_{1})$ suppose namely $w^{\prime}$ is inside the region $a_{1}yw^{\prime\prime}y^{\prime}$ then replace $w^{\prime}$ by $a_{1}$, which takes us to another case. So we must have $y\in N(a_{1})$ and $y^{\prime}\in N(a_{2})$ in this case since the vertices of $O^{\prime}$ separate $C$ into two regions, if $w$ is in the region of $a_{1}$ replace it by $a_{1}$, otherwise replace $w$ and the vertex of $O^{\prime}$ in the region of $a_{2}$ by $y$ and $a_{2}$. Other cases can be treated similarly.
Reduction Rule 8
Suppose $I\neq\emptyset$ and $O^{\prime}=\emptyset$. Color any vertex $u$ in $C$ blue except if one of the following is true:
  • $u\in I^{\prime}$
  • $\exists u^{\prime}\in N(y)$ where $y\in N(u)\cap Y$ and $\{u,u^{\prime}\}$ dominates $I$.
  • $Y\subset N(a_{1})$ and $\exists u^{\prime}\in N(Y)$ where $\{u,u^{\prime}\}$ dominates $I-N(a_{1})$
Soundness. Suppose $O\textquoteright=\emptyset$ ’ = ∅, let u be a vertex that is not blue, $u\notin I\textquoteright$ ’. Then by reduction rule 6 rule $N(u)\cap Y=\{y\}$ for some $y\in Y$. There must be some $u\textquoteright\in I\cup B\cup\{a1,a2\}$ which is in $S$. Let $W$ be as in the above reduction rule and we have that no four vertices of $W$ can be in $S$. If $Y\subset N(a_{1})$ then if $a_{1}\in W$, $u$ and $u^{\prime}$ dominate $I-N(a1)$ and $u^{\prime}\notin P$ so $u^{\prime}\in N(Y)$ by reduction rule 6, so $u$ follows one of the above cases. Else if $|W|=3$ replace the three vertices of $W$ by $a_{1}$, $a_{2}$ and a vertex of $Y$ having 2 or more neighbors in $W$ so $|W|=2$ and we are also done since $u^{\prime}\in N(y)$ otherwise replace $u$ and $u^{\prime}$ by $a_{1}$ and $a_{2}$. Now if $Y\not\subset N(a_{1})$ suppose $|W|=3$ then as before we can always replace the vertices of $W$ including u, so $|W|=2$ and $u^{\prime}\in N(y)$ otherwise $u$ and $u^{\prime}$ can be replaced.
Reduction rules $9-13$ concern blue vertices. They are applicable on general graphs and should be applied successively and exhaustively.
Reduction Rule 9
Let $v\in V(G)$ where $d(v)\leq 1$. If $\exists w$ where $N_{h}[v]\subset N(w)$ and there exists at most one vertex $z\in N(v)-w$ with $d(z)\geq 2$, then color $v$ blue.
Soundness. Suppose $v\in S$. If $w\notin S$, replace $v$ by $w$, otherwise replace $v$ by $z$.
Reduction Rule 10
Delete all the edges between blue vertices, and delete 0-vertices that are blue.
Reduction Rule 11
Let $v$ be a blue vertex, if $N(v)=\{u,w\}$ for some $u,w\in V(G)$ and the edge $uw$ exists, delete it and decrease each of $d(u)$ and $d(w)$ by one.
Soundness. $v$ is a blue vertex so it is a high demand vertex due to reduction rule 10, thus one of its neighbors must be in $S$. Suppose (namely) $u$ is in $S$, then in $G-S$ the edge $uw$ is deleted and $w$ must have $d(w)-1$ other neighbors in $S$, so we can replace $d(w)$ by $d(w)-1$, and $u$ would not need any other neighbor in $S$ so decreasing $d(u)$ has no effect on the solution.
Reduction Rule 12
Let $v$ be a blue vertex. If for some 1-vertex $u\neq v$ $N(v)\subset N[u]$, then replace $d(u)$ by 0.
Soundness. $v$ is blue which means it is a high demand vertex by reduction rule 10. Thus $N(v)\cap S\neq\emptyset$ so $N(u)\cap S\neq\emptyset$ regardless of $d(u)$, so it will make no difference to decrease $d(u)$ to zero.
Reduction Rule 13
Let $u$ be a 0-vertex where $|N(u)|=2$. If there exists a 0-vertex $v$ where $N(u)=N(v)$, then delete $u$.
Soundness. Suppose $u\in S$, then if $v\notin S$ replace $u$ by $v$, otherwise replace $u$ and $v$ by $N(u)$. Thus $u\notin S$ and since $d(u)=0$ we can delete it.
Lemma 4
After applying the reduction rules above to a graph any candidate region $C$ contains at most $15$ vertices.
The above lemma follows by the planarity of the graph and the above reduction rules. Below is the worst case example for the region $C$ when $O^{\prime}\neq\emptyset$. In this case $w$ dominates $I$ with a vertex of $O^{\prime}$ and the same applies to $u$. (see Figure 2)
[…]
Figure 2: Worst case when $O^{\prime}\neq\emptyset$
Theorem 3.1
Let $G$ be a planar graph, then computing a vector dominating set for $G$ of size at most $k$ can be reduced to computing a vector dominating set of size at most $k$ to a planar graph $G^{\prime}$ where $|G^{\prime}|\leq 101k$.
Proof
We will show that after applying the above reduction rules we obtain $|G^{\prime}|\leq 101k$. Suppose $G$ admits a solution $S$ with a maximal $S$-region decomposition $\mathscr{R}$. Let $x$, $y$ and $z$ be the numbers of type 1, 2 and 3 outer paths of regions in $\mathscr{R}$ respectively. Let $A^{\prime}=A-V(\mathscr{R})$, $L^{\prime}=L-V(\mathscr{R})$, $A_{1}=\{c\in A^{\prime}:|N(c)|\geq 3\}$ and $A_{2}=A^{\prime}\setminus A_{1}$. Let $K=\{v:d(v)=1$ and $v$ belongs to some outer path of }$\mathscr{R}\}\cup S$. The number of outer paths of $\mathscr{R}$ is at most $6k$ (since we have at most $3k$ regions), with two 1-vertices at most on each outer path, thus $|K|\leq 12k+k=13k$.
Claim1: Each vertex of $A_{1}\cup L$ has three or more neighbors in $K$. To see this, take $c\in A_{1}$ then $N(c)\cap L=\emptyset$ (a leech cannot be adjacent to a 0-vertex) so $\forall v\in N(c)$, $v\in\mathscr{R}$ thus $v$ is on some outer path by planarity and since $|N(c)|\geq 3$ Claim1 follows. Now take $l\in L^{\prime}$. $l$ is a leech for some type 3 path, so $\exists s,v\in N(l)$ with $s\in S$ and $d(v)=1$. $v$ must be on some outer path since $v\in V(\mathscr{R})$ ($v\notin A\cup L)$ and $v$ cannot be in the interior of a region due to planarity. Now, by reduction rule 11, $l$ must have a neighbor $v^{\prime}$ different from $s$ and $v$. We must have $d(v^{\prime})=1$ since $l$ is a leech and $v^{\prime}\notin L$ since vertices of $L$ are colored blue by reduction rule 9, so no two leeches can be adjacent. Thus similar to $v$, $v^{\prime}$ is on some outer path. This proves our claim.
By considering the planar bipartite graph formed of the vertices of $A_{1}\cup L$ and $K$ and the edges between them, the number of edges $e$ is bounded above by $42(|A_{1}|+|L|+|K|)-$ 4, thus we have $43(|A_{1}|+|L|)\leq e\leq 2(|A_{1}|+|L|+13k)-$ 4 which gives $|A_{1}|+|L|\leq 26k$.
Claim2. Let $c\in A_{2}$, then any vertex $v\in N(c)$ is on an type 1 or 3 outer path of a region $R\in\mathscr{R}$. This is true since $v$ is on a type 1 path $p$ by definition of $A$. If $p$ is an outer path, then Claim2 follows, otherwise $p$ must cross some region $R\in\mathscr{R}$ and $v$ must be on the boundary of $R$ by planarity. But then this cannot occur if the outer path containing $v$ is of type 2, which proves Claim2.
Consequently (due to Claim2), the number of vertices in $N(A_{2})$ is at most $x+z$. Now no two vertices of $A_{2}$ can have the same neighborhood by reduction rule 13, so by viewing each vertex of $A_{2}$ as an edge between vertices of $N(A_{2})$, we can say that $|A_{2}|\leq 3(x+z)-6$.
Thus we have $|V(G^{\prime})|\leq 15\cdot 3k+(x+3y+2z)+(3x+3z)+26k=71k+4x+3y+5z$ but this expression is maximized if $z=6k$, which gives $|V(G^{\prime})|\leq 101k$. ∎